Mar is 68 years old and john is 44. How many years ago was marvin age twice johns age?
Answer by checkley77(12569)

68-x=2(44-x)
68-x=88-2x
-x+2x=88-68
x=20 ans. years ago Marvan was twice John's age.
Proof:
68-20=2(44-20)
48=2*24
48=48